OpenStudy (anonymous):

In a geometric sequence, a(down)7=14 and r = – 1/2. What is the 10th term of the sequence?

OpenStudy (dls):

\[\LARGE a_7=14 \]and \[\LARGE r=\frac{-1}{2}\] for \[\LARGE a_{10}=ar^{9}\] can you find a from given conditions? \[\LARGE a_7=ar^{n-1}\]

OpenStudy (dls):

you are given a7=14 and r=-1/2 and n=6

OpenStudy (anonymous):

yes

OpenStudy (dls):

a=?

OpenStudy (anonymous):

I'm confused on how to fill it out correctly

OpenStudy (dls):

\[\LARGE 14=a(\frac{1}{2})^6\]

OpenStudy (dls):

calculate a

OpenStudy (anonymous):

ok 1 sec

OpenStudy (dls):

its wrong,sorry

OpenStudy (dls):

\[\LARGE a=14 \times (2)^6\]

OpenStudy (dls):

relax urself :) take your time!

OpenStudy (anonymous):

I'm confused

OpenStudy (dls):

yes :)

OpenStudy (dls):

a=896

OpenStudy (dls):

now substitute here.. \[\LARGE a_{10}=ar^{9}\] a=896 r=-1/2

OpenStudy (anonymous):

I keep getting the wrong answer :(

OpenStudy (dls):

what did you get?

OpenStudy (anonymous):

\[x_{10}=-\frac{ 1 }{ 512 }\]

OpenStudy (dls):

?????how

OpenStudy (anonymous):

can you show me your work

OpenStudy (dls):

\[\LARGE a_{10}=896 \times \frac{-1}{512}\]

OpenStudy (anonymous):

so it's b. -1.75

OpenStudy (anonymous):

The geometric sequence is \[a_n = ar^{n-1}\] Given: \[r = -\frac{1}{2}\] \[a_7 = ar^6 =14\] Substituting r into ar^6 gives: \[a\left(-\frac{1}{2}\right)^6 = 14\] \[a = 14 * 2^6 = 14 * 64 = 896\] Substiting a and r into the definition of a geometric sequence gives us the sequence we're after: \[a_n = 896\left(-\frac{1}{2}\right)^{n-1}\] so \[a_{10} = 896\left(-\frac{1}{2}\right)^9 = -\frac{896}{512} = -1\frac{3}{4} \]